A kettle draws the power 2000 W and has an efficiency of 85%. You fill the boiler with 1.5 liters of water with a temperature of 15 ° C

A) How long does it take at least to boil the water in the kettle?
B) The kettle's thermostat broke and the water boils even after it has reached its boiling point. Determine how long it takes to boil all the water from the kettle
1
Expert's answer
2021-03-14T19:17:14-0400

a)


"\eta Pt=Q=mc(T_2-T_1)\\(0.85)(2000)t=(1.5)(4200)(100-15)\\
t=315\ s=5.25\ min"

b)


"\eta Pt=Q=mc(T_2-T_1)+mR\\(0.85)(2000)t=\\(1.5)(4200)(100-15)+1.5(2256\cdot10^3)\\
t=2306\ s=38.4\ min"